OpenStudy (anonymous):

Solve by elimination. equation 1;3x-2y=16 & equation 2;6x+2y=20

OpenStudy (mertsj):

Add the two equations.

OpenStudy (anonymous):

add the two equations and you will see that the y's cancel each other out. Now you can find x. Once x is found, sub that back into either of the original equations and solve for y.

OpenStudy (anonymous):

3x-2y=16 + 6x+2y=20 = (9x+36)/9 x=4 6(4)+2y=20 24+2y=20 -24 -24 = (2y=-4)/2 y=-2 3(4)-2(-2)=16 12-(-4)=16 12+4=16 16=16 6(4)-2(-2)=20 24+(-4)=20 24-4=20

OpenStudy (anonymous):

worked it out for ya

OpenStudy (anonymous):

Thank You very Much .:))